HELP PLEASE

Most automobiles have a coolant reservoir to catch
Answer #1

\DeltaT = change in temperature = 95 - 13 = 82

V = initial Volume = 21.5 L = 21.5 x 0.001 m3 = 0.0215 m3

\DeltaV = change in Volume

Change in Volume of radiator is given as

\DeltaVrad = (0.0215) (51 x 10-6) (82) = 0.0000899

Change in Volume of fluid is given as

\DeltaVfluid = (0.0215) (400 x 10-6) (82)

Volume spilled over :

\DeltaVfluid - \DeltaVrad = (0.0215) (400 x 10-6) (82) - (0.0215) (51 x 10-6) (82) = 6.15 x 10-4 m3